# Urban Region Pre-training and Prompting: A Graph-based Approach

### v1

@@ A.4. Appendix of attribute graph kernel function
Extending the above definition to the scenario of graphs with node attributes, we use $\mathbf{X}\in\mathbb{R}^{n\times d}$ to denote the attribute matrix of a graph with $n$ nodes and attribute dimension $d$. Specifically, the node attribute matrices of $G_{1},G_{2}$ are denoted by $\mathbf{X}_{1}\in\mathbb{R}^{n_{1}\times d},\mathbf{X}_{2}\in\mathbb{R}^{n_{2}\times d}$. Thus, for the graph $G_{\times}=G_{1}\times G_{2}$, the node attribute matrix is $\mathbf{S}=\mathbf{X}_{1}\mathbf{X}_{2}^{\top},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$, where the item $(i,j)$ encodes the similarity between the $i$-th node in $G_{1}$ and the $j$-th node in $G_{2}$. We flattened $\mathbf{S}$ to $\mathbf{s}\in\mathbb{R}^{n_{1}n_{2}}$ to obtain the formula for the random walk kernel:
(3) $K_{p}(G_{1},G_{2})=\sum_{i,j=1}^{|V_{\times}|}\mathbf{s}_{i}\mathbf{s}_{j}[A_{\times}^{p}]_{ij}=\mathbf{s}^{\top}A^{p}_{\times}\mathbf{s}.$
A.5. Appendix of dataset statistics and sources
We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI). The region division is based on census tracts and the urban imagery are 256 x 256 pixels with a resolution of 4.7 meters. The detailed statistics and sources of the dataset is shown in Table 5.
@@ A.5. Appendix of dataset statistics and sources
Imagery 7,160 - ArcGIS (Arcgis, 2024), UUKG (Liu et al., 2023c)
@@
Check-in 106,902 - Foursquare (Yang et al., 2014; Foursquare, 2024)
@@
A.6. Appendix of baselines description
We compare our model with representative baseline models, including graph embedding methods, knowledge graph embedding methods and state-of-the-art region embedding methods.
@@ I. Knowledge graph embedding methods:
  • TransR (Lin et al., 2015) captures the complex interactions between entities and relationships in the knowledge graph by projecting them into different spaces and learning their representation in these spaces.
  • TransR-N method uses the embedding of the region node learned by TransR as the region representation.
  • TransR-G method is to average the embedding of the region subgraph learned by TransR as the region representation.
@@ II. Graph embedding methods:
  • node2vec (Grover and Leskovec, 2016) generates a sequence of nodes by performing random walks on the graph to capture the neighborhood information of nodes, and then learns the representation of nodes using the Skip-Gram model.
  • GAE (Kipf and Welling, 2016) consists of an encoder and a decoder. The encoder is used to learn the representation of nodes, and the decoder tries to reconstruct the adjacency matrix from these node representation.
@@ III. State-of-the-art urban region embedding methods:
  • MVURE (Zhang et al., 2021) constructs intra- and inter-region data multi-view based on human mobility and region attributes, and then performs multi-view fusion to learn region representation.
  • MGFN (Wu et al., 2022) builds human mobility patterns and then learns intra- and inter-pattern relationships to obtain region representation through a multi-level cross-attention mechanism.
  • HREP (Zhou et al., 2023a) designs a multi-relational heterogeneous graph of regions and used a relation-aware GCN to learn region representation, as well as a prefix prompt mechanism for downstream tasks.
  • HRE is a part of HREP without prompt learning.
  • ReCP (Li et al., 2024) is a multi-view comparative prediction model that learns region representation by capturing the unique information of a single view as well as the consistency between two views.
A.7. Appendix of Implementation
In our experiments, all the region embedding dimensions is set to 144. The number of HGT layer and head is 2 and 4, and the triplet margin parameter $\delta=2$, fusion loss weight $\mu=0.01$. We set the learning rate to 0.001 and the weight decay to 1e-6. In the task-learnable prompt module, we set the number of random walk kernel to 2 and the size of prompt graph to 6 and 8 of each layer.
We implemented our model with Pytorch 1.8, and conducted experiments on a machine equipped with a NVIDIA RTX 3090 GPU, a 64 GB RAM and a 3.50 GHz CPU.
A.8. Appendix of prompt graphs
Figure 10 displays additional learned prompt graphs, with the majority of them being circular structural features.
Figure 10. Visualization of additional learned prompt graphs.

### v4

@@ A.4. Attribute Graph Kernel Function
Extending the above definition to the scenario of graphs with node attributes, we use $\mathbf{X}\in\mathbb{R}^{n\times d}$ to denote the attribute matrix of a graph with $n$ nodes and attribute dimension $d$. The node attribute matrices of $G_{1},G_{2}$ are denoted by $\mathbf{X}_{1}\in\mathbb{R}^{n_{1}\times d},\mathbf{X}_{2}\in\mathbb{R}^{n_{2}\times d}$. Thus, for the graph $G_{\times}=G_{1}\times G_{2}$, the node attribute matrix is $\mathbf{S}=\mathbf{X}_{1}\mathbf{X}_{2}^{\top},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$, where the item $(i,j)$ encodes the similarity between the $i$-th node in $G_{1}$ and the $j$-th node in $G_{2}$. We flattened $\mathbf{S}$ to $\mathbf{s}\in\mathbb{R}^{n_{1}n_{2}}$ to obtain the formula for the random walk kernel:
$K_{p}(G_{1},G_{2})=\sum_{i,j=1}^{|V_{\times}|}\mathbf{s}_{i}\mathbf{s}_{j}[A_{\times}^{p}]_{ij}=\mathbf{s}^{\top}A^{p}_{\times}\mathbf{s}.$
A.5. Experiment Details
A.5.1. Dataset Statistics and Sources.
Statistics and sources of the datasets are shown in Table 6. The region division is based on census tracts and the urban imagery are 256 x 256 pixels with a resolution of 4.7 meters. For carbon emissions data provided in 1x1 $km$ grid cells, we map them to our regional divisions. If a region covers more than one grid cell, the emissions from all relevant cells are summed. For the income and education data, each region is assigned a specific value representing the average income or number of people with an education for that region.
@@ A.5. Experiment Details
Imagery 7,160 - ArcGIS (Arcgis, 2024; Liu et al., 2023d)
@@
Check-in 106,902 - Foursquare (Foursquare, 2024)
@@
Carbon 1x1 $km$ 1x1 $km$ ODIAC (Oda and Maksyutov, 2015)
Income 180 - NYCOD (NYC OpenData, 2024)
Education - 869 CHIDP (Chicago Data Portal, 2024)
A.5.2. Baselines Descriptions.
Here we provide more details about the three types of baseline methods.
@@ I. Knowledge graph embedding methods:
  • $\circ$
    TransR (Lin et al., 2015) projects entities and relationships into relation-specific spaces. TransR-N uses the region node embedding as the representation. TransR-G averages embeddings of all nodes in the region subgraph.
@@ II. Graph embedding methods:
  • $\circ$
    node2vec (Grover and Leskovec, 2016) uses random walks to generate node sequences and learns representations with the Skip-Gram model.
  • $\circ$
    GAE (Kipf and Welling, 2016) has an encoder for node representations and a decoder for reconstructing the adjacency matrix.
@@ III. State-of-the-art urban region embedding methods:
  • $\circ$
    MVURE (Zhang et al., 2021) constructs multi-view data based on human mobility and region attributes for region representation.
  • $\circ$
    MGFN (Wu et al., 2022) learns relationships among human mobility patterns through cross-attention.
  • $\circ$
    HREP (Zhou et al., 2023a) creates a multi-relational graph and a relation-aware GCN with a non-graph prefix prompt. HRE is a variant without the prompt module.
  • $\circ$
    ReCP (Li et al., 2024a) learns region representation through a multi-view fusion that captures unique and consistent information.
A.5.3. Implementation.
We use five-fold cross-validation to evaluate GURPP and baselines on multiple tasks across both cities. In GURPP, the region embedding dimension is set to 144 (Zhou et al., 2023a). The HGT layer and head numbers are 2 and 4, respectively; triplet margin $\delta=2$; fusion loss weight $\mu=0.01$; learning rate is 0.001; and weight decay is 1e-6. In the task-learnable prompt module, we set the random walk kernel layers to 2, prompt graph sizes to 6 and 8, and region subgraph node size to 50. Due to the lack of satellite image datasets for CHI, we removed the image encoding module and adjusted the fusion layer dimensions accordingly.
Our model is implemented with PyTorch 1.8 and DGL 2.2.1+cu118 on a machine with an NVIDIA RTX 3090 GPU, 64 GB RAM, and a 3.50 GHz CPU. Scalability experiments use an NVIDIA H100 80GB GPU, 1 TB RAM, and a 3.2 GHz CPU.
A.5.4. Comparison with Language Prompt.
We conduct a comparison with LLMs, including Deepseek R1, Gemini 2 Flash, in predicting the top 10 regions with the highest crime probabilities in NYC. The specific variant of GURPP used in this comparison is $\text{GURPP}_{M}$-P4. We provide each language model with detailed descriptions of the 180 regions in NYC, including region IDs and information about POIs. The prompt is as follows.
I will provide information on the division of 180 regions in Manhattan, New York, with each region including a region ID and detailed description. Please analyze the potential crime risk of each area based on their descriptions, rank them according to the likelihood of criminal incidents from highest to lowest, and list the top 10 highest-risk regions. The regional information is formatted as: {mhtr0: mhtr0 is located in Manhattan, encompassing 1318 POIs, with the main types being Restaurants and Other Eating Places, Clothing Stores …} …